OpenStudy (anonymous):

A shelf in the Metro Department Store contains 95 colored ink cartridges for a popular ink-jet printer. Six of the cartridges are defective. a) customer selects 2 cartridges at random, what is the probability that both are defective? (round answer to 5 decimal places) b) customer selects 2 cartridges at random, what is the probability that at least 1 is defective? (round answer to 3 decimal places)

OpenStudy (anonymous):

I'm pretty sure A) is (6/95)(5/94)=.00335946 and it's saying the answer is wrong when rounded

OpenStudy (anonymous):

B) 1-(answer to A). Need help!

OpenStudy (kropot72):

The answer to a) is given by \[\frac{6\times5}{95\times94}\] What was your result when rounded?

OpenStudy (anonymous):

I figured out A, but B isn't the right equation I don't think

OpenStudy (kropot72):

b) The answer is not found by subtracting the answer to a) from 1. The solution can be found from: P(at least one defective) = 1 - P(0 defective) \[P(0\ defective)=\frac{89\times88}{95\times94}=you\ can\ calculate\]

OpenStudy (anonymous):

got it, thanks!

OpenStudy (kropot72):

Good. What was your result?